# Deep Regression Representation Learning with Topology

### v3

@@ H.2 Age estimation on AgeDB-DIR dataset
H.3 Super-resolutuion on DIV2K dataset
We exploit the DIV2K dataset (Timofte et al., 2017) for 4x super-resolution training (without the 2x pretrained model) and we evaluate on the validation set of DIV2K and the standard benchmarks: Set5 (Bevilacqua et al., 2012), Set14 (Zeyde et al., 2012), BSD100 (Martin et al., 2001), Urban100 (Huang et al., 2015). We follow the setting of Lim et al. (2017) and exploit their small-size EDSR model as our baseline architecture. We adopt the standard metric PNSR and SSIM. The trade-off parameters $\lambda_{d}$ and $\lambda_{t}$ are set to $0.1$ and $1$, respectively. Table 3 shows that both ${\mathcal{L}}_{d}$ and ${\mathcal{L}}_{t}$ contribute to improving the baseline and adding both terms has the largest impact.
@@ H.5 Different improvement gap between synthetic and real-world datasets
However, on NYU-Depth-V2, the representation of the head part (with many samples in the target space) is relatively well learned, leading to a smaller impact on our regularizers. This might be a reason for the lower improvement. From the regression baseline (Figure 2(a)), the representation in the blue region (head part) already shows a lower intrinsic dimension and collapses like a line so there is little opportunity for our regularizers to have a strong impact (and hence the smaller improvement on the MAE (Many)). In contrast, the impact on the synthetic dataset and the green region (corresponding to the tail part, with limited samples in the target space) is more significant, resulting in a higher improvement.

### v4

@@ H.2 Age estimation on AgeDB-DIR dataset
H.3 Super-resolution on DIV2K dataset
We use the DIV2K dataset (Timofte et al., 2017) for 4x super-resolution training (without the 2x pretrained model) and we evaluate on the validation set of DIV2K and the standard benchmarks: Set5 (Bevilacqua et al., 2012), Set14 (Zeyde et al., 2012), BSD100 (Martin et al., 2001), Urban100 (Huang et al., 2015). We follow the setting of Lim et al. (2017) and exploit their small-size EDSR model as our baseline architecture. We adopt the standard metrics PNSR and SSIM. The trade-off parameters $\lambda_{d}$ and $\lambda_{t}$ are set to $0.1$ and $1$, respectively. Table 3 shows that both ${\mathcal{L}}_{d}$ and ${\mathcal{L}}_{t}$ contribute to improving the baseline and adding both terms has the largest impact.
@@ H.5 Different improvement gap between synthetic and real-world datasets
However, on NYU-Depth-V2, the representation of the head part (with many samples in the target space) is relatively well-learned, leading to a smaller impact on our regularizers. This might be a reason for the lower improvement. From the regression baseline (Figure 2(a)), the representation in the blue region (head part) already shows a lower intrinsic dimension and collapses like a line so there is little opportunity for our regularizers to have a strong impact (and hence the smaller improvement on the MAE (Many)). In contrast, the impact on the synthetic dataset and the green region (corresponding to the tail part, with limited samples in the target space) is more significant, resulting in a higher improvement.